OpenStudy (anonymous):

Which equation is related to this equation? b + 17 = 52 A. b = 52 + 17 B. b = 52 ∙ 17 C. b = 52 ÷ 17 D. b = 52 – 17

OpenStudy (bishoppatton):

d.

OpenStudy (calculusxy):

When you will be solving for b, you need to isolate it. That means that you need to keep by itself. Regularly, if you were to solve for b, what would you do first?

OpenStudy (anonymous):

are u sure bishop

OpenStudy (bishoppatton):

yes.... 35 + 17 = 52

OpenStudy (anonymous):

Use a related equation that has the variable on one side. Then simplify the other side. 18b = 18 A. b = 0 B. b = 36 C. b = 324 D. b = 1

OpenStudy (bishoppatton):

This is pretty simple... you should know this

OpenStudy (anonymous):

d

OpenStudy (bishoppatton):

good job

OpenStudy (bishoppatton):

lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Use a related equation that has the variable on one side. Then simplify the other side. A. B. C. D.

OpenStudy (anonymous):

wha lol

OpenStudy (bishoppatton):

?

OpenStudy (bishoppatton):

3

OpenStudy (bishoppatton):

6 divided by 3 equals 2

OpenStudy (anonymous):

12 divided by 6 =d right

OpenStudy (anonymous):

or 2

OpenStudy (bishoppatton):

yes

OpenStudy (bishoppatton):

so the blank would be 12

OpenStudy (anonymous):

Use a related equation that has the variable on one side. Then simplify the other side. p + 6 = 18 A. p = 3 B. p = 12 C. p = 24 D. p = 108

OpenStudy (bishoppatton):

solve by using 18-6

OpenStudy (bishoppatton):

18-6=?

OpenStudy (anonymous):

12

OpenStudy (bishoppatton):

the answer is p = 12

OpenStudy (anonymous):

we got a 100000000000 woooooohoooooo

OpenStudy (bishoppatton):

CONGRATS!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

im gonna give u 20 medals

OpenStudy (bishoppatton):

How could i get 20 medals?

OpenStudy (anonymous):

ill close u join next and ill give u medals

OpenStudy (bishoppatton):

ok... i get it